Please help me solve the following:
Which ordered pair is the solution of the following system of equations?
3x+2y=4
-2x + 2y= 24
A. (-4,8) B. (-4,-8) C. (2,-5) D. (2,-1)
How do you get this answer
Thank you,
Natasha

Answer by Alan3354(69443)

3x + 2y=4
-2x +2y= 24
-------------- Subtract
5x + 0y = -20
x = -4
------
Sub for x in either eqn, find y.
(-4,8)